OpenStudy (anonymous):

Solve the system of equations. x + 3y = −1 2x + 2y = 6

OpenStudy (susanna):

x + 3y = −1 2x + 2y = 6 Use elimination -2x-6y=2 2x+2y=6 -2x and 2x cancel each other out. -4y=8 y=-2

OpenStudy (anonymous):

ohh that makes sense

OpenStudy (susanna):

plug in y=-2 into one of the equations, and solve for x. x+3(-2)=-1 x-6=-1 x=5

OpenStudy (susanna):

check (5, -2) 2(5) + 2(-2)= 6 10-4=6 6=6

OpenStudy (anonymous):

thank you so much!! :)

OpenStudy (susanna):

sure